OpenStudy (anonymous):

Marylou is blowing up a beach ball. The package says that the ball has a diameter of 18 inches. What is the volume of the beach ball when it is completely full? Use 3.14 for . 24,416.6400000000 cubic inches 3,052.08 cubic inches 339.12 cubic inches 1,017.36 cubic inches

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@perl

OpenStudy (aum):

Volume of a sphere V = \(\Large \frac 43 \pi r^3; ~~~ r = 9 \text{ in.}\)